1800Pa 6500Pa 3600Pa 2700Pa 5100Pa
Answer #1

I got \(\mathrm{B}, 6500\) Pascals.

First I used \(\rho=M / V\), so:

\(917=M /\left(.75^{3}\right)\)

\(\mathrm{M}=386.8594 \mathrm{~kg} .\)

Then \(I\) used \(P=F / A\) \(P=\frac{F}{A}=\frac{\operatorname{mgeos}(16)}{.75_{D}^{2}}=6478.8558Pa\)